# A sixteen multiple-amplifier-sensing CCD and characterization techniques targeting the next generation of astronomical instruments

### v1

@@ Abstract
Gate removal efficiency (GRE), 16 multiple-amplifier sensing CCD (MAS-CCD), correlated noise analysis, nondestructive readout sensor, single-electron resolution imager, single-photon counting imager
@@ 1 Introduction
Low-noise silicon imagers have been identified as a key technology for the construction of the next generation of scientific spectroscopic experiments. The low light signal projected on the pixels due to the spectral dispersion on faint objects could be obscured by the uncertainty added by different background sources. In particular, for terrestrial observations, the readout noise of the sensor could be a considerable contribution in the blue region of the spectrum where the sky background is suppressed. For example, this regime is particularly relevant for the planned Stage-5 Spectroscopic Survey Experiment (Spec-S5) [1], which would measure $\sim 10^{8}$ distant galaxies (an order of magnitude more than current surveys) for redshift from 2 to 5 to study the mechanism driving the expansion of the universe after inflation.
@@
Recent studies in [2] show that the signal-to-noise ratio of this spectral line can be increased using the non-destructive readout of Skipper Charge Coupled Devices (Skipper-CCD) to reduce the readout noise contribution in the pixel measurement. At the same time, preliminary studies presented as part of the Snowmass process in [3, 4] show that the redshift successful measurement of LGB galaxies can be increased by approximately 25% by reducing the readout noise down to 1e- compared to the RMS noise level of around 3e- in the Dark Energy Survey Instrument [5]. This opens an opportunity to increase the survey speed of future terrestrial spectroscopic surveys by improving the readout noise of the detectors.
@@
Another spectroscopic application seeking low-noise silicon sensor technology is the search for Earth-like planets in the habitable zones of Sun-like stars using a coronagraph instrument in space [6]. In this case, the visible and near-infrared bands contain an abundance of information about exoplanet atmospheres. The low expected flux, in the order of a few photons per hour per pixel [7], requires a sub-electron noise for single-photon detection and fast readout (total exposure time of around 60 seconds) to avoid excessive occupancy of cosmic ray traces in the sensor. For this case, the non-destructive readout of the Skipper-CCD[8], has been identified as a candidate solution.
@@
The Skipper-CCD provides a powerful way, by multiple measurements of the collected charge, to reduce the readout noise of the pixel at the expense of an extra read time due to the multiple sampling that is not affordable for this kind of application. An extended version of it, called the Multiple-Amplifier-Sensing Charge Coupled Device (MAS-CCD), was recently presented in [9] and its principles demonstrated in [10] provides a solution to overcome the extra read time of Skipper-CCDs. The multiple inline architecture of the MAS-CCD, as shown in the simplified schematic of Fig. 1a provides an interesting solution to meet this requirement without increasing the total read time compared to the current devices used on the current experiments [5]. The pixel charge packet is measured sequentially by the amplifiers and the final pixel value is computed with the use of the multiple non-destructive measurements taken. The proof of concept of this new architecture has been recently presented for a sensor with eight amplifiers in the output stage [10]. This paper extends these results to a detector of sixteen output stages which allows reaching the noise operation regime expected for Spec-S5. At the same time, the paper provides a new theoretical framework, techniques, and tools to characterize and optimize the new features of the sensor compared to regular CCDs. In particular, the article provides the method to optimally mix the information from the multiple amplifiers and gives a model for a new source of charge transfer inefficiency, which we call gate removal inefficiency, related to the extraction from the sense node back to the serial register pixels.
@@
In the following section, an introduction to MAS architecture is addressed. In Sec. 3 digital signal processing and optimum averaging techniques are presented to further improve the detector performance. A mathematical framework to model Gate Removal Efficiency in MAS detectors is addressed in Sec. 4. Experimental results are shown in Sec. 5. Finally, concluding remarks are given in Sec. 6.
@@ 2 MAS-CCD readout technique and noise performance analysis
Figure 1a shows a MAS simplified schematic, its basic components are the pixel matrix, a bent serial register, and many inline amplifiers. In a basic readout scheme the pixel charge packets are transferred from the matrix to the serial register, and then transferred and readout across the multi amplifiers stages (A1 to A16).
@@
Figure 1b presents a simplified view of the components and gates of the readout stages at the end of the serial register of the sensor. For each stage, there is the sense node (SN) connected to the gate of the transistors (Ai) to measure the charge in the serial register channel. The pixel separation gate (PS) is used to remove the charge from the SN after its readout to put it under the horizontal gates (H1, H2, and H3). The three gates form the structure for one pixel in the serial register, where the rest of the charge is stored. A dump gate is located at the end of the last amplifier for charge disposal. For more detailed description refer to [9, 10]
@@
where $s_{i,j}$ is the $j$ sample from the amplifier $i$, and $N_{s}$ and $N_{a}$ are the total number of samples taken by each amplifier and the total number of amplifiers, respectively. The readout uncertainty contribution is reduced as the square root of the number of samples $N_{s}$ as demonstrated for the Skipper-CCD [8, 11, 12]. With the MAS-CCD, the noise of each amplifier can be modeled as $n_{i}=n_{Ai}+n_{CM}$ with $n_{Ai}$ the independent noise of each amplifier and $n_{CM}$ the common noise source affecting all the amplifier stages, as shown in Fig. 1b. As explained in the following section, there are tools that can be used to monitor and reduce the effect of this correlated noise in the final readout noise performance of the sensor. For the calculations presented in this article the pixel values $s_{i,j}$ are computed using the dual-slope integrator (DSI) processing technique [13].
When $n_{CM}\ll n_{Ai}$ (or $n_{CM}=0$) the MAS readout noise is reduced by the square root of $N_{a}$ [10]. When the noise from the different amplifiers is independent but not necessarily equal, the noise uncertainty in the final pixel value is
$\sigma_{ind}=\dfrac{\sqrt{\sum_{i=1}^{N_{a}}\sigma_{i}^{2}}}{N_{a}}.$ (2)
@@
where $\sigma^{2}_{n_{CM}}$ is the variance contribution common mode noise at the output of the MAS. $\sigma^{2}_{n_{CM}}$ can be evaluated using the power spectrum of the common noise in the system and the response of the MAS-CCD readout operation. The frequency response of the MAS-CCD is calculated in Appendix A and compared to the frequency responses of the single and multiple DSI of the Skipper-CCD.
@@ 3.1 Gain equalization
where $s_{k}$ is the sample value of a charged pixel from the $k$-th amplifier. $c_{i}$ measures the ratio of the gain of the $i$ amplifier relative to the average gain of all the amplifiers. The pixel value $s_{k}$ must be large enough to be distinguishable from readout and background noise, and small enough to avoid saturated values.
@@ 3.2 Optimum average of the channels
$\text{min}_{A}\ A^{T}\Sigma A,\ A^{T}\mathbf{1}=1$
where $\mathbf{1}$ represents a column vector of ones and dimensions $N_{a}\times 1$. Weights are then obtained solving the problem for the vector of weights $A$, resulting in
@@ 3.3 Noise decorrelation
where if $i=1$, every element starting at $1$ is instead considered starting at $2$. The square matrix represents the covariance matrix of the channels containing empty pixels, and the column vector on the right-hand side represents the covariances of the $i$-th channel with respect to the rest of the channels.
A potential use of this type of technique is in spectroscopy where projected spectral lines appear separated by a few pixels on the CCD surface . If the intermediate pixels have a low background contribution, their pixel value can be used to decorrelate the noise from channels measuring actual spectra. This could add a link between the length of pixel separation between amplifiers and the space between projected spectral lines in the active region.
4 Gate removal efficiency (GRE) of the sense nodes charge packets
@@
For the model, we assume that a fraction of the charge is left behind in the sense node of each amplifier when the packet is taken out to be transferred to the next amplifier. To differentiate this phenomenon from the standard CTI/CTE process in CCDs, the efficiency for removing the charge from the sense nodes is defined as gate removal efficiency (GRE) and gate removal inefficiency (GRI). This mechanism can be explained using Fig. 1b. Once the charge in the channel of the serial register is measured in each output stage by the amplifiers (Ai) through the non-destructive sense nodes (SN), the charge should be moved out using the Pixel Separation gate (PS). The PS voltage is moved down with H1. Then, the typical three-phase clock sequence allocates the charge under H2 before the next pixel readout. With this sequence, the charge is removed from the SN and incorporated into the next serial register pixels. During the charge removal from the SN, some of the carriers could be left behind. A fraction of the measured charge by one amplifier could stay in the sensor and be added to the following pixel charge carrier. One of the critical aspects is that this extra charge does not affect the measured value of the next pixel since it will be part of the reference voltage for the DSI calculation (the pedestal level), however, this fraction of extra charge will affect the next amplifier. Looking at the output images, one of the characteristic signatures of this effect is that the first amplifier in the chain does not see the effect of the GRE, even in a situation where the effect is very aggressive in the other channels.
@@
The GRE process can be thought as depicted in the block diagram of Fig. 2a. The first amplifier (stage 1) measures the charge without GRE issues, illustrated by the meter before the GRE process. After that the GRE process takes place, distorting the charge packet. The next amplifier stage will measure this effect in the next time instant (represented by the ideal delay block $z^{-1}$).
4.1 GRE between two consecutive amplifiers
The “GRE process” block in Fig. 2a can be modeled as a recursive discrete system. Following the methodology in [15, Chapter 3, pp.103-108] we write the equation for the value of the next charge packet $q_{i}[n+1]$ (time instant $n+1$) measured in amplifier $i$ as a function of the previous charge packet $q_{i}[n]$ (time instant $n$) measured in the same amplifier and the charge packet measured in the predecessor amplifier $q_{i-1}[n]$.
@@ 4.1 GRE between two consecutive amplifiers
where $(1-\epsilon)$ is the GRE and $\epsilon$ the GRI.
@@
It is worth noting that solving Eq.(10) for the GRI results in
@@
models the distortion effect of the GRE process on the charge packet measurement between two consecutive amplifiers.
A block diagram of this process is shown in Fig. 2b, where for each stage, the pixel value is obtained first, the GRE process is modeled by the distortion transfer function $D(z)$ and the pure delay $z^{-1}$ models the deferred measurement between two consecutive inline amplifiers. In other words, if the GRE was perfect ($\epsilon=0$) then $Q_{i}(z)=z^{-1}Q_{i-1}(z)$ and since $z^{-1}$ is just a pure delay, anti-transforming results in $q_{i}[n]=q_{i-1}[n-1]$, i.e., the charge in amplifier $i$ in time instant $n$ is exactly equal to the charge measured in the previous amplifier $i-1$ in the previous time instant $n-1$.
4.2 GRE in the $N_{a}$ inline amplifiers
From the point of view of the GRE, the distortion effect in the $N_{a}$ inline amplifiers, can be seen as a cascade connection of the $D(z)$ transfer functions. An important point to make is that the first amplifier measures the charge packet without GRI related issues because in the first sense node no charge was lost yet. Therefore, this measurement can be considered as the ideal input signal to the rest of $N_{a}-1$ inline amplifiers. To model the effect of the GRE at any of the amplifiers, after the first one, the cascading of distortion transfer functions results in
@@ 4.2 GRE in the $N_{a}$ inline amplifiers
with $2\leq n_{a}\leq N_{a}$. This transfer function can be used to predict the effect of the GRE for a given stream of input charge packets. A well-known input could be obtained by applying a flat field of light to the sensor followed by a readout with the $N_{a}$ amplifiers that extend beyond the number of columns of the active region, producing overscan pixels that should be empty if GRE is perfect. This is similar to the extended pixel edge response (EPER) method for CTI [13] and results in a negative step-function input of charge for the $N_{a}$ inline amplifiers.
Assuming as input to $D_{n_{a}}(z)$ a negative step function of $Q$ electrons of charge, with z-transform $-Q/(1-z^{-1})$, the output is the product of the two transfer functions: $-QD_{n_{a}}(z)/(1-z^{-1})$. Anti-transforming by partial fraction expansions of the $n_{a}-1$ order pole in 13 and subtracting the result to the total charge $Q$ gives the output $y_{n_{a}}[n]$ for the $n_{a}$ amplifier caused by GRE due to a negative step of charge:
@@
were $u[n]$ is the unit step function, $2\leq n_{a}\leq N_{a}$ and $n\geq 0$. Equation (14) predicts the output in the overscan pixels ($n=0$ is the first overscan pixel) for each of the amplifiers.
@@
$y_{1}[n]=Q\epsilon^{n+1}u[n]$ (15)
which is an exponential response. Usually, for CTI measurements with EPER method, only the first overscan pixel is measured. For this amplifier this results in $y_{1}[0]=Q\epsilon$, which as expected corresponds to the original charge packet $Q$ times the GRI, the inefficiency for removing the charge.
@@
4.3 Average GRE/GRI of the amplifier chain
In the previous description, $\epsilon$ represents the individual GRI for each amplifier. In the MAS-CCD presented in this work, the final pixel value is obtained by averaging the measurement in the $N_{a}$ amplifiers. Therefore, the real impact on the final pixel value of the gate removal efficiency modeled in this section should be computed as the average of the outputs in each of the amplifiers. Assuming a relatively low $\epsilon$ most of the impact of the GRI is present in the first overscan pixel $y_{n_{a}}[0]$, using Eq.(16), the average GRI for the amplifier chain can be computed by averaging $y_{n_{a}}[0]$ for $2\leq n_{a}\leq N_{a}$ and dividing by the total charge $Q$

### v2

@@ Abstract
Node removal efficiency (NRE), 16 multiple-amplifier sensing CCD (MAS-CCD), correlated noise analysis, nondestructive readout sensor, single-electron resolution imager, single-photon counting imager
@@ 1 Introduction
Low-noise silicon imagers have been identified as a key technology for the construction of the next generation of scientific spectroscopic experiments. The low light signal projected on the pixels due to the spectral dispersion on faint objects could be obscured by the uncertainty added by different background sources. In particular, for terrestrial observations, the readout noise of the sensor could be a considerable contribution in the blue region of the spectrum where the sky background is suppressed. For example, this regime is particularly relevant for the planned Stage-5 Spectroscopic Survey Experiment (Spec-S5) [1], which would measure $\sim 10^{8}$ distant galaxies (an order of magnitude more than current surveys) for redshift from 2 to 5 times the original wavelength to study the mechanism driving the expansion of the universe after inflation.
@@
Recent studies in [2] show that the signal-to-noise ratio of this spectral line can be increased using the non-destructive readout of Skipper Charge Coupled Devices (Skipper-CCD) to reduce the readout noise contribution in the pixel measurement. At the same time, preliminary studies presented as part of the particle physics community future planning in the US (known as the Snowmass process) in [3, 4] show that the rate of the successful measurement of the redshift of Lyman-Break galaxies can be increased by approximately 25% by reducing the readout noise down to 1e- compared to the RMS noise level of around 3e- in the Dark Energy Survey Instrument [5]. This opens an opportunity to increase the survey speed of future terrestrial spectroscopic surveys by improving the readout noise of the detectors.
@@
Another spectroscopic application seeking low-noise silicon sensor technology is the search for Earth-like planets in the habitable zones of Sun-like stars using a coronagraph instrument in space [6]. In this case, the visible and near-infrared bands contain abundant information about exoplanet atmospheres. The low expected flux, in the order of a few photons per hour per pixel [7], requires sub-electron noise for single-photon detection and fast readout (total exposure time of around 60 seconds) to avoid excessive occupancy of cosmic ray traces in the sensor. For this case, the non-destructive readout of the Skipper-CCD[8], has been identified as a candidate solution.
@@
The Skipper-CCD provides a powerful way, by multiple measurements of the collected charge, to reduce the readout noise of the pixel at the expense of an extra read time due to the multiple sampling which is not tolerated for this kind of application. An extended version of it, called the Multiple-Amplifier-Sensing Charge Coupled Device (MAS-CCD), was recently presented in [9] and its principles demonstrated in [10] provides a solution to overcome the extra read time of Skipper-CCDs. Its good performance encouraged further characterization efforts presented in [11, 12, 13]. The multiple inline architecture of the MAS-CCD, as shown in the simplified schematic of Fig. 1a, measures the pixel charge packet sequentially, and the final pixel value is computed using the multiple non-destructive measurements taken. This provides an interesting solution to meet the requirement without increasing the total read time compared to the current devices used in the current experiments [5]. This paper extends the results in 5 to a detector of sixteen output stages which allows reaching the noise operation regime expected for Spec-S5. At the same time, the paper provides a new theoretical framework, techniques, and tools to characterize and optimize the new features of the sensor compared to regular CCDs. In particular, the article provides the method to optimally mix the information from the multiple amplifiers and gives a model for a new source of charge transfer inefficiency, which we call node removal inefficiency, related to the extraction from the sense node back to the serial register pixels.
@@
In the following section, an introduction to MAS architecture is addressed. In Sec. 3 digital signal processing and optimum averaging techniques are presented to further improve the detector performance. A mathematical framework to model Node Removal Efficiency in MAS detectors is addressed in Sec. 4. Experimental results are shown in Sec. 5. Finally, concluding remarks are given in Sec. 6.
@@ 2 MAS-CCD readout technique and noise performance analysis
Figure 1a shows a MAS-CCD simplified schematic, its basic components are the pixel matrix, a bent serial register, and 16 inline amplifiers. In a basic readout scheme, the pixel charge packets are transferred from the matrix to the serial register, and then transferred and readout across the multi amplifiers stages (A1 to A16).
@@
Figure 1b presents a simplified view of the components and gates of the readout stages at the end of the serial register of the sensor. For each stage, there is a sense node (SN) connected to the gate of the transistors (Ai) to measure the charge in the serial register channel. The pixel separation gate (PS) is used to remove the charge from the SN after its readout to put it under the horizontal gates (H1, H2, and H3). These three gates form the structure for one pixel in the serial register, where the rest of the charge is stored. A dump gate is located at the end of the last amplifier for charge disposal, for a more detailed description refer to [9, 10].
@@
where $s_{i,j}$ is the $j$ sample from the amplifier $i$, and $N_{s}$ and $N_{a}$ are the total number of samples taken by each amplifier and the total number of amplifiers, respectively. The readout uncertainty contribution is reduced as the square root of the number of samples $N_{s}$ as demonstrated for the Skipper-CCD [8, 14, 15]. With the MAS-CCD, the noise of each amplifier can be modeled as $n_{i}=n_{Ai}+n_{CM}$ with $n_{Ai}$ the independent noise of each amplifier and $n_{CM}$ the common noise source affecting all the amplifier stages, as shown in Fig. 1b. As explained in the following section, there are tools that can be used to monitor and reduce the effect of this correlated noise in the final readout noise performance of the sensor. For the calculations presented in this article the pixel values $s_{i,j}$ are computed using the dual-slope integrator (DSI) processing technique [16, Chapter 6, pp.537-541].
When $n_{CM}\ll n_{Ai}$ (or $n_{CM}=0$) the MAS readout noise is also reduced by the square root of $N_{a}$ [10]. When the noise from the different amplifiers is independent but not necessarily equal, the noise uncertainty in the final pixel value is
$\sigma_{ind}=\dfrac{\sqrt{\sum_{i=1}^{N_{a}}\sigma_{i}^{2}}}{N_{a}}$ (2)
@@
where $\sigma^{2}_{n_{CM}}$ is the variance contribution common mode noise at the output of the MAS. $\sigma^{2}_{n_{CM}}$ can be evaluated using the power spectrum of the common noise in the system and the frequency response of the MAS-CCD readout operation. The frequency response of the MAS-CCD is calculated in Appendix A and compared to the frequency responses of the single and multiple DSI of the Skipper-CCD.
@@ 3.1 Gain equalization
where $s_{k}$ is the sample value of a charged pixel from the $k$-th amplifier. $c_{i}$ measures the ratio of the gain of the $i$ amplifier relative to the average gain of all the amplifiers. The pixel value $s_{k}$ must be much larger than the total uncertainty ($s_{k}>>\sigma_{pix}$), but small enough to avoid saturated values.
@@ 3.2 Optimum average of the channels
$\text{min}_{A}\ A^{T}\Sigma A,\text{subject to}\ A^{T}\mathbf{1}=1$
where $\mathbf{1}$ represents a column vector of ones and dimensions $N_{a}\times 1$ and the condition is set so that the sum of the $\alpha_{i}$ coefficients is 1. Weights are then obtained solving the problem for the vector of weights $A$, resulting in
@@
For example, solving this equation to find a generic solution for $A^{T}=[\alpha_{1},\alpha_{2}]$ (assuming a $\Sigma$ matrix of order 2 ($N_{a}=2$)) leads to $\alpha_{1}=(\sigma_{2,2}^{2}-\sigma_{c}^{2})/(\sigma_{2,2}^{2}+\sigma_{1,1}^{2}-2\sigma_{c}^{2})$ and $\alpha_{2}=(\sigma_{1,1}^{2}-\sigma_{c}^{2})/(\sigma_{2,2}^{2}+\sigma_{1,1}^{2}-2\sigma_{c}^{2})$, where $\sigma_{c}=\sigma_{1,2}=\sigma_{2,1}$. From here, it is possible to evaluate different conditions. If $\sigma_{c}^{2}\ll\sigma_{2,2}^{2}$ and $\sigma_{c}^{2}\ll\sigma_{1,1}^{2}$, therefore $\alpha_{1}\approx\sigma_{2,2}^{2}/(\sigma_{2,2}^{2}+\sigma_{1,1}^{2})$ and $\alpha_{2}\approx\sigma_{1,1}^{2}/(\sigma_{2,2}^{2}+\sigma_{1,1}^{2})$.
Considering equal contributions of noise $\sigma_{2,2}^{2}=\sigma_{1,1}^{2}$ leads to $\alpha_{1}=\alpha_{2}=1/2$. Another option is to consider different contribution of noise $\sigma_{2,2}^{2}=3\sigma_{1,1}^{2}$ gives $\alpha_{1}=3/4$ and $\alpha_{2}=1/4$. This reflects coefficients are inversely proportional to the channel noise contribution.
@@ 3.3 Noise decorrelation
where if $i=1$, every element starting at $1$ is instead considered starting at $2$. The square matrix represents the covariance matrix of the channels containing empty pixels, and the column vector on the right-hand side represents the covariances of the $i$-th channel with respect to the rest of the channels. This technique is applied before the optimum average of the channels, presented in Sec.3.2.
A potential use of this type of technique is in spectroscopy where projected spectral lines appear separated by a few empty pixels on the CCD[18]. If the intermediate pixels have a low background contribution, their pixel value can be used to decorrelate the noise from channels measuring actual spectral lines. This could add a link between the length of pixel separation between amplifiers and the space between projected spectral lines in the active region.
4 Node removal efficiency (NRE) of the sense nodes charge packets
@@
For the model, we assume that a fraction of the charge is left behind in the sense node of each amplifier when the packet is taken out to be transferred to the next amplifier. To differentiate this phenomenon from the standard CTI/CTE process in CCDs, the efficiency for removing the charge from the sense nodes is defined as node removal efficiency (NRE) and node removal inefficiency (NRI). This mechanism can be explained using Fig. 1b. Once the charge in the channel of the serial register is measured in each output stage by the amplifiers (Ai) through the non-destructive sense nodes (SN), the charge should be moved out using the Pixel Separation gate (PS). The PS voltage is moved with H1. Then, the typical three-phase clock sequence allocates the charge under H2 before the next pixel readout. With this sequence, the charge is removed from the SN and incorporated into the next serial register pixels. During the charge removal from the SN, some of the carriers could be left behind. A fraction of the measured charge by one amplifier could stay in the sensor and be added to the following pixel charge carrier. One of the critical aspects is that this extra charge does not affect the measured value of the next pixel since it will be part of the reference voltage for the DSI calculation (the pedestal level), however, this fraction of extra charge will affect the next amplifier. Looking at the output images, one of the characteristic signatures of this effect is that the first amplifier in the chain does not see the effect of the NRE, even in a situation where the effect is very aggressive in the other channels.
@@
The NRE process can be thought as depicted in the block diagram of Fig. 2a. The first amplifier (stage 1) measures the charge without NRE issues, illustrated by the meter before the NRE process. After that the NRE process takes place, distorting the charge packet. The next amplifier stage will measure this effect in the next time instant (represented by the ideal delay block $z^{-1}$).
4.1 NRE between two consecutive amplifiers
The “NRE process” block in Fig. 2a can be modeled as a recursive discrete system. Following the methodology in [19, Chapter 3, pp.103-108] we write the equation for the value of the next charge packet $q_{i}[n+1]$ (time instant $n+1$) measured in amplifier $i$ as a function of the previous charge packet $q_{i}[n]$ (time instant $n$) measured in the same amplifier and the charge packet measured in the predecessor amplifier $q_{i-1}[n]$.
@@ 4.1 NRE between two consecutive amplifiers
where $(1-\epsilon)$ is the NRE and $0\leq\epsilon\leq 1$ the NRI.
@@
It is worth noting that solving Eq.(10) for the NRI results in
@@
models the distortion effect of the NRE process on the charge packet measurement between two consecutive amplifiers.
A block diagram of this process is shown in Fig. 2b, where for each stage, the pixel value is obtained first, the NRE process is modeled by the distortion transfer function $D(z)$ and the pure delay $z^{-1}$ models the deferred measurement between two consecutive inline amplifiers. In other words, if the NRE was perfect ($\epsilon=0$) then $Q_{i}(z)=z^{-1}Q_{i-1}(z)$ and since $z^{-1}$ is just a pure delay, anti-transforming results in $q_{i}[n]=q_{i-1}[n-1]$, i.e., the charge in amplifier $i$ in time instant $n$ is exactly equal to the charge measured in the previous amplifier $i-1$ in the previous time instant $n-1$.
4.2 NRE in the $N_{a}$ inline amplifiers
From the point of view of the NRE, the distortion effect in the $N_{a}$ inline amplifiers, can be seen as a cascade connection of the $D(z)$ transfer functions. An important point to make is that the first amplifier measures the charge packet without NRI related issues because in the first sense node no charge was lost yet. Therefore, this measurement can be considered as the ideal input signal to the rest of $N_{a}-1$ inline amplifiers. To model the effect of the NRE at any of the amplifiers, after the first one, the cascading of distortion transfer functions results in
@@ 4.2 NRE in the $N_{a}$ inline amplifiers
with $2\leq n_{a}\leq N_{a}$. This transfer function can be used to predict the effect of the NRE for a given stream of input charge packets. A well-known input could be obtained by applying a flat field of light to the sensor followed by a readout with the $N_{a}$ amplifiers that extend beyond the number of columns of the active region, producing overscan pixels that should be empty if NRE is perfect. This is similar to the extended pixel edge response (EPER) method for CTI [16, Chapter 5, pp.423-429] and results in a negative step-function input of charge for the $N_{a}$ inline amplifiers.
Assuming a negative step function of $Q$ electrons (i.e., the charge decreases from 0 to $-Q$) with z-transform $-Q/(1-z^{-1})$, as input to $D_{n_{a}}(z)$, the output in the z-domain is given by $Y_{-Q}(z)=-QD_{n_{a}}(z)/(1-z^{-1})$. The time-domain output $y_{n_{a}}[n]$ for the $n_{a}$-th amplifier, caused by the NRE due to a negative step of charge (from $Q$ to 0 electrons), is $y_{n_{a}}[n]=\mathcal{Z}^{-1}\{Y_{-Q}(z)\}+Q$ for $n\geq 0$. The anti-transform is performed by partial fraction expansion for the pole of order $n_{a}-1$ in (13), obtaining:
@@
where $u[n]$ is the unit step function, $2\leq n_{a}\leq N_{a}$ and $n\geq 0$. Equation (14) predicts the output in the overscan pixels ($n=0$ is the first overscan pixel) for each of the amplifiers.
@@
$y_{2}[n]=Q\epsilon^{n+1}u[n],$ (15)
which is an exponential response. Usually, for CTI measurements with EPER method, only the first overscan pixel is measured. For this amplifier this results in $y_{2}[0]=Q\epsilon$, which as expected corresponds to the original charge packet $Q$ times the NRI, the inefficiency for removing the charge.
@@
4.3 Average NRE/NRI of the amplifier chain
In the previous description, $\epsilon$ represents the individual NRI for each amplifier. In the MAS-CCD presented in this work, the final pixel value is obtained by averaging the measurement in the $N_{a}$ amplifiers. Therefore, the real impact on the final pixel value of the node removal efficiency modeled in this section should be computed as the average of the outputs in each of the amplifiers. Assuming a relatively low $\epsilon$ most of the impact of the NRI is present in the first overscan pixel $y_{n_{a}}[0]$, using Eq.(16), the average NRI for the amplifier chain can be computed by averaging $y_{n_{a}}[0]$ for $2\leq n_{a}\leq N_{a}$ and dividing by the total charge $Q$